OpenStudy (anonymous):

Frogs have been breeding like flies at the Enormous State University (ESU) campus! Each year, the pledge class of the Epsilon Delta fraternity is instructed to tag all the frogs residing on the ESU campus. Two years ago (t = 0) they managed to tag all 54,000 of them (with little Epsilon Delta Fraternity tags). This year's pledge class discovered that all the tags had all fallen off, and they wound up tagging a total of 86,400 frogs. (a) Find an exponential model for the frog population.

OpenStudy (mertsj):

86400=54000e^k(2)

OpenStudy (mertsj):

Solve for k

OpenStudy (anonymous):

so i got 1.6=e^k(2)

OpenStudy (anonymous):

then ln1.6=e^k(2)(ln)

OpenStudy (anonymous):

does ln and e cancel out?

OpenStudy (mertsj):

\[\ln 1.6=2k \ln e\]

OpenStudy (anonymous):

ln 1.6 = 2k

OpenStudy (anonymous):

1.6/2 = k

OpenStudy (mertsj):

They don't cancel out.

OpenStudy (anonymous):

ln e = 1

OpenStudy (mertsj):

\[\ln _{e}e=1\]

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

.08 = k

OpenStudy (mertsj):

\[\frac{\ln 1.6}{2}=k\]

OpenStudy (mertsj):

k=.235

OpenStudy (anonymous):

ooo forgot about ln

OpenStudy (mertsj):

\[\ln 1.6 = 2k\]

OpenStudy (anonymous):

what do i do with k=.235 , can i plug that in so i can make the equation now?

OpenStudy (anonymous):

y=Ab^x

OpenStudy (mertsj):

\[A=a _{0}e ^{.235t}\]

OpenStudy (anonymous):

i plugged in f(t)=54000e^.235t

OpenStudy (anonymous):

but its not correct

OpenStudy (mertsj):

Then leave out the 54000

OpenStudy (anonymous):

so just f(t)=e^.235t?

OpenStudy (mertsj):

Did you see what I wrote?

OpenStudy (mertsj):

A=a0e.235t

OpenStudy (mertsj):

Or it is possible that they want more or fewer significant digits on the constant.

OpenStudy (anonymous):

is the zero supposed to be in the equation you showed me?

OpenStudy (mertsj):

\[A _{o}\] indicates the initial amount. Use anything you please for that.

OpenStudy (mertsj):

Probably you should use the same variables that your book or instructor has been using.